# Cardinality of Real Numbers and Set Axiomatization

### v2

@@ 2 Discussion
The properties of the set $S^{\prime}$ clearly demonstrate that it supports the Löwenheim-Skolem theorem rather than the diagonal argument. This result, along with the Löwenheim-Skolem theorem, shows that our understanding of the nondenumerability of real numbers consists in our knowing what it is for this to be proved and not in our grasp of a model.
@@
The set $S^{\prime}$, while complete and countable, does not provide an explicit sequence of all real numbers. This means that for a specific given number, it is not possible to identify its particular $n$-th position in the sequence. This limitation arises because the numbers are generated arbitrarily by the function $f$. For any given $n$-th position, $f$ may assign infinitely many numbers within its interval of operation. To address this, Example 2 considers a set equivalent to $S^{\prime}$, which is also generated by inserting transcendental numbers between algebraic numbers. Contrary to the previous example, this method specifies all generated transcendental numbers at each step, thus providing a sequence in which numbers have clearly defined positions. This approach merges the concepts of Gödel’s Constructible Universe (V = L) with the earlier proposed Wang’s $\Sigma$ model [19], which differs from ZF models. However, instead of using Wang’s original work [19], let us modify the model introduced in [20].
@@
The model begins with a 00-th layer, which consists of some denumerable totality of objects, which may be taken to be, for instance, the positive integers or all finite sets built up from the empty set. In our next example, we will consider that the 00-th layer comprises algebraic numbers. Then, the first layer includes all the objects of the 00-th layer and, in addition, all those sets of these objects that correspond to conditions that contain bound variables ranging over objects of the 00-th layer, but no bound variables ranging over the first or higher layers. Generally, the $n+1$-st layer contains all the objects of the $n$-th layer, along with all such sets of these objects determined by conditions whose bound variables range over objects of the $n$-th layer at most. The hierarchy of layers is continued beyond the finite ordinals. Layer $\omega$, for instance, is the sum of all finite layers, and layer $\omega+1$ contains, in addition, also such sets as are determined by conditions whose bound variables range over the objects of layer $\omega$ at most. Since all the sets of layer $\alpha$ are enumerable by a function $E_{\alpha}$ ranging over entities of a layer, all sets of any subsystem $\Sigma_{\alpha}$ of $\Sigma$ are therefore enumerable. Therefore, this model requires that any infinite set can be enumerated in an appropriate partial system of $\Sigma$. We consider here Wang’s model, but the conception of cumulative layers (or orders) is not original with Wang [22, 23], and there are also other attempts to realize it [21].
@@
In the next example, we present our model. Its 00-th layer consists of a set of algebraic numbers. Layer one includes all numbers from the 00-th layer and all transcendental numbers created by simultaneously applying an infinite number of diagonal constructions on the 00-th layer. The layer $n+1$ is obtained by applying $\omega$ diagonal constructions on layer $n$, which also incorporates all previous layers. More details are provided below.
@@
Example 2. Let us consider an interval of numbers $[M,W]$, starting by writing a denumerable sequence of all algebraic reals $A_{\nu}$ that belong to that interval, as specified in the corresponding array (12). Each number will be represented by an infinite binary sequence, even if it has a finite number of digits, with each digit $a_{\nu\mu}$ being either 0 or 1:
@@
The sequence $S_{r}$ contains only all algebraic numbers, and each algebraic number has a well-defined place within that sequence. Let us now use Cantor’s diagonal construction to generate numbers not contained in this sequence. Such numbers are defined by sequences of digits $d_{\nu 1},d_{\nu 2},\ldots,d_{\nu\mu},\ldots$, where each $d_{\nu\mu}$ is either 0 or 1, determined such that $d_{\nu\nu}\neq a_{\nu\nu}$. From this sequence of $d_{\nu\mu}$, we formulate the element $D_{1}=(d_{11},d_{22},\ldots,d_{\nu\nu},\ldots)$, ensuring that $D_{1}\neq A_{\nu}$ for any index $\nu$. Thus, a number different from all numbers in the sequence containing all algebraic numbers is created. Therefore, it must be a transcendental number. This will be the first transcendental number generated and will be incorporated into the sequence $S_{r}$ as the first element, which shifts all elements in the original sequence from the $n$-th to the $n+1$-th place. This recursive construction will be repeated; it will be applied again on the newly generated set $S_{r}$, generating in the same way the second transcendental number and placing it in the sequence $S_{r}$ in the first position, shifting all elements in $S_{r}$ again, then the third transcendental number, and so on. Please note that all numbers generated in this way are well-defined. All digits $d_{\nu\mu}$ of such generated transcendental numbers $D_{1},D_{2},\ldots,D_{\nu},\ldots$ are defined and known. This is because all digits in the initial sequence $S_{r}$ that contains all algebraic reals are defined. In creating transcendental numbers by Cantor’s diagonal construction, there is no freedom because each diagonal element 0 must be switched to 1, or 1 to 0; one must apply $d_{\nu\nu}\neq a_{\nu\nu}$ for each diagonal element. Since in the initial sequence $S_{r}$ of algebraic numbers all $a_{\nu\mu}$ are known and fixed, all numbers $D_{1}$, $D_{2}$, …that are produced with the diagonal construction are defined.
@@
To generate the first layer in our model, the construction will be modified so that in the first step, not only one but infinitely many transcendental numbers are generated. This is possible because we can use all decimal places $a_{11},a_{12},\ldots,a_{1\mu},\ldots$ simultaneously and apply an infinite number of diagonal constructions simultaneously in the first step. There will be a diagonal construction that starts with the first decimal place $a_{11}$, one that starts with the second decimal place $a_{12}$, one with the third $a_{13}$, and so on. New transcendental numbers will be created using an infinite number of diagonal constructions, each starting from a different decimal place. After step one, the first layer, denoted as $\Sigma_{1}$, will consist of an infinite sequence of $\aleph_{0}$ algebraic numbers from the 00-th layer and an infinite sequence of $\aleph_{0}$ transcendental numbers produced by an infinite number of diagonal constructions applied to the set of algebraic reals of the 00-th layer. These two sets, the 00-th layer and the transcendental numbers generated by diagonal constructions, will be combined into one sequence by alternating the first number from the 00-th layer with the first number from the list of created transcendental numbers, and so on, to create the first layer, the $\Sigma_{1}$ set. The construction from step one will be repeated, applying an infinite number of diagonal constructions to the sequence that represents the just created set $\Sigma_{1}$, creating an additional $\aleph_{0}$ transcendental numbers that will then be combined with set $\Sigma_{1}$ to create set $\Sigma_{2}$. This recursive construction will continue indefinitely, creating an additional $\aleph_{0}$ of newly created transcendental numbers each time.
@@
However, this construction cannot continue indefinitely; it will stop after $\omega$ steps due to the atomistic structure of real numbers, which prevents endless division of an interval. The process stops exactly after $\omega$ divisions because an infinite division of intervals results in singletons that cannot be divided further. At each step of creating layers, infinite amounts of transcendental numbers are inserted between the algebraic numbers of the 00-th layer, dividing each segment defined by a pair of algebraic numbers further until it becomes a singleton, as follows from the BW theorem or the NIP.
@@
In this way, this example is equivalent to Example 1. In both cases, transcendental numbers are generated and inserted between algebraic numbers, further subdividing the intervals into smaller and smaller segments. The model also represents a variant of forcing, and the forcing technique can be implemented here. Additionally, the diagonal argument cannot be used to create missing numbers not generated by this model because it is already utilized to generate new numbers. Any ‘missing’number produced by diagonal construction is, in fact, just a number that belongs to this model.
@@
Remarks to the non-denumerability proofs
There are many variations of proofs of the uncountability of real numbers. However, fundamentally, all of them share a common basis: they revolve around infinite nested sequences of intervals. We will discuss the critical point that appears in all these proofs, the interpretation of the intersection of these intervals and the nature of the number that resides within this intersection. According to the NIP, the intersection of an infinite nested sequence of closed nonempty intervals must be a zero-size interval, a singleton, or a single number. This number must be one of the elements from the intervals of the sequence, which, in all uncountability proofs, are the numbers of the supposed denumerable sequence assumed to include all numbers. Contrarily, in all these proofs, that number is declared as the number that is outside the denumerable sequence, and its existence is used as the argument that the sequence does not contain all numbers. This discrepancy between the NIP, which posits that the number is a member of the sequence, and the uncountability proofs, which assert the opposite, will be elaborated upon, considering the proofs from 1873 and 1879, and the diagonal argument.
@@
In the concluding arguments of both proofs, two fundamental issues arise: the technical problem that the proposed construction of a number different from all listed numbers cannot be completed, and the logical problem that assumes if a number does not have an assigned specific $n$-th place in the sequence, it is not in the sequence. It will be demonstrated that the number proposed to be different from all others cannot be constructed as intended. What is actually constructed is the number $\eta$, which is not distinct from other numbers. The construction requires an additional step to differentiate it from the listed numbers, but this final step cannot be completed as proposed. Also, the constructed number $\eta$ does not have a specific $n$-th place in the sequence due to the method of its construction; however, this does not imply that it is not part of the sequence. Let us elaborate on both of these points.
@@
The claim is that in the limit as $\nu\to\infty$, where $\omega_{a}^{\nu}=A$ and $\omega_{b}^{\nu}=B$, and when $A=B$, there exists an open interval $(A,B)$ that eliminates all endpoints $\omega_{a}^{\nu}$, $\omega_{b}^{\nu}$, and is not empty, containing the number $A=B$. To quote: “ because $\eta$ is in the interior of the interval $[\omega_{a}^{\nu},\omega_{b}^{\nu}]$, but $\omega_{\nu}$ lies outside of it. For all $\nu,\eta\in(\omega_{a}^{\nu},\omega_{b}^{\nu})$ but $\omega_{\nu}\notin(\omega_{a}^{\nu},\omega_{b}^{\nu})$.”
@@
The numbers $\eta$ and $\omega_{n}$ are then defined as the intersections of an infinite sequence of nested intervals
@@
In the nested interval representation, the equivalent expression stipulates that at each finite step $n$, the process involves considering the intersection (17) of the first $n$ intervals $I_{n}$ representing the first $n$ decimals of the numbers $\eta$, and $\omega_{n}$. Within these intersections, distinct subintervals $H_{n}$ for $\eta$ and $\Omega_{n}$ for $\omega_{n}$ are selected to ensure divergence. This divergence is created by choosing disjunctive subintervals to guarantee that $\eta$ and $\omega_{n}$ differ at the $n$-th stage, effectively making the numbers different at each $n$-th decimal place or $n$-th interval, because they will have different intersections (17):
@@
The diagonal argument assumes that the process of creating disjunctive subintervals $H_{i}$ and $\Omega_{i}$ can continue indefinitely, theoretically allowing for the construction of the number $\eta$ different from all listed numbers $\omega_{n}$. However, this seemingly straightforward assumption contradicts the NIP. At each finite stage in sequence (16), every interval $I_{n}=[\alpha_{n},\beta_{n}]$ has a finite size, with $\alpha_{n}<\beta_{n}$ consistently satisfied. This ensures that each interval $I_{n}$ contains at least two disjunctive subintervals $H_{n}$ and $\Omega_{n}$, necessary to construct number $\eta$ such that it will be different from listed numbers $\omega_{n}$. However, this intuitively natural requirement cannot be met when applied to the entire sequence (16) as its intersection interval (17), in the limit, has zero size. In the limit, as $n\rightarrow\infty$, $\alpha_{n}\rightarrow A$ and $\beta_{n}\rightarrow B$, and $A=B$, the intersection (17) of all intervals $I_{n}=[\alpha_{n},\beta_{n}]$, i.e., the intersection interval $[A,B]$, has zero size and contains only a single number. So, there is no possibility that in that intersection interval will exist two disjunctive subintervals $H_{n}$ and $\Omega_{n}$, which is required by diagonal argument to make the intersection (18) representing number $\eta$ different from the intersections (19) representing a listed number $\omega_{n}$. The intersection theorem dictates that sequence (16) has a zero-size intersection with only one number in it. Intervals $H_{n}$ representing decimals $a_{n}$ in number $\eta$ and intervals $\Omega_{n}$ representing $\omega_{nn}$ in $\omega_{n}$ for every $n$ belong to the same interval $I_{n}=[\alpha_{n},\beta_{n}]$ because they represent the same $n$-th decimal position. Therefore, they must lie within the same intersection (17) of intervals $I_{n}$ at each step for any $n$. For every $n$ in the construction of numbers $\eta$ and $\omega_{n}$, the intersection of the first $n$ intervals $[\alpha_{n},\beta_{n}]$ has non-zero size and contains disjunctive subintervals $H_{n}$ and $\Omega_{n}$, allowing $\eta$ to differ from the first $n$ listed numbers $\omega_{n}$ in their $n$-th decimal places. However, at these finite stages, number $\eta$ is not yet defined or constructed, regardless of how large $n$ is. Number $\eta$ is only defined or constructed in the limit as $n\rightarrow\infty$, when $\alpha_{n}\rightarrow A$ and $\beta_{n}\rightarrow B$. But in this limit, the intersection $[A,B]$ is a singleton containing only one number. The requirement at each $n$-th step to select two disjunctive subintervals $H_{n}\neq\Omega_{n}$ that make sequence intersections (18) and (19) different, cannot be fulfilled in the limit because the intersection (17) becomes a singleton, a single number. So, the number $\eta$, that is different from all listed numbers cannot be constructed since in the limit do not exist two disjunctive subintervals, required to make the number $\eta$ different from all listed $\omega_{n}$ numbers.

### v3

@@ 4 Explicit sequence of everywhere dense set of real numbers
The properties of the set $S_{m}$ clearly demonstrate that it supports the Löwenheim-Skolem theorem rather than the diagonal argument. This result, along with the Löwenheim-Skolem theorem, shows that our understanding of the nondenumerability of real numbers consists in our knowing what it is for this to be proved and not in our grasp of a model.
@@
The set $S_{m}$, while complete and countable, does not provide an explicit sequence of all real numbers. This means that for a specific given number, it is not possible to identify its particular $n$-th position in the sequence. This limitation arises because the intervals in the set $S_{m}$ even countable are not enumerated and to each interval is assigned one irrational number that is not specified. To address this, considered will be a set equivalent to $S_{m}$, which is generated by inserting irrational numbers between rational numbers. Contrary to the previous example, this method specifies all generated irrational numbers at each step, thus providing a sequence in which numbers have clearly defined positions. This approach merges the concepts of Gödel’s Constructible Universe (V = L) with the earlier proposed Wang’s $\Sigma$ model [17], which differs from ZF models. However, instead of using Wang’s original work [17], let us modify the model introduced in [18].
@@
The model begins with a 00-th layer, which consists of some denumerable totality of objects, which may be taken to be, for instance, the positive integers or all finite sets built up from the empty set. In our next theorem, we will consider that the 00-th layer comprises rational numbers. Then, the first layer includes all the objects of the 00-th layer and, in addition, all those sets of these objects that correspond to conditions that contain bound variables ranging over objects of the 00-th layer, but no bound variables ranging over the first or higher layers. Generally, the $n+1$-st layer contains all the objects of the $n$-th layer, along with all such sets of these objects determined by conditions whose bound variables range over objects of the $n$-th layer at most. The hierarchy of layers is continued beyond the finite ordinals. Layer $\omega$, for instance, is the sum of all finite layers, and layer $\omega+1$ contains, in addition, also such sets as are determined by conditions whose bound variables range over the objects of layer $\omega$ at most. Since all the sets of layer $\alpha$ are enumerable by a function $E_{\alpha}$ ranging over entities of a layer, all sets of any subsystem $\Sigma_{\alpha}$ of $\Sigma$ are therefore enumerable. Therefore, this model requires that any infinite set can be enumerated in an appropriate partial system of $\Sigma$. We consider here Wang’s model, but the conception of cumulative layers (or orders) is not original with Wang [19, 20], and there are also other attempts to realize it [21].
@@
In the next theorem, we present our model. Its 00-th layer consists of a set of rational numbers. Layer one includes all numbers from the 00-th layer and all irrational numbers created by simultaneously applying an infinite number of diagonal constructions on the 00-th layer. The layer $n+1$ is obtained by applying $\omega$ diagonal constructions on layer $n$, which also incorporates all previous layers. More details are provided below.
Theorem 4 (Explicit sequence of everywhere dense and complete set of real numbers).
The construction described below generates an explicit sequence that contains all real numbers.
@@
Let us consider an interval of numbers $[M,W]$, starting by writing a denumerable sequence of all rational reals $A_{\nu}$ that belong to that interval, as specified in the corresponding array (15). Each number will be represented by an infinite binary sequence, even if it has a finite number of digits, with each digit $a_{\nu\mu}$ being either 0 or 1:
@@
The sequence $S_{r}$ contains only all rational numbers, and each rational number has a well-defined place within that sequence. Let us now use Cantor’s diagonal construction to generate numbers not contained in this sequence. Such numbers are defined by sequences of digits $d_{\nu 1},d_{\nu 2},\ldots,d_{\nu\mu},\ldots$, where each $d_{\nu\mu}$ is either 0 or 1, determined such that $d_{\mu\mu}\neq a_{\nu\nu}$. From this sequence of $d_{\nu\mu}$, we formulate the element $D_{1}=(d_{11},d_{12},\ldots,d_{\nu\mu},\ldots)$, ensuring that $D_{1}\neq A_{\nu}$ for any index $\nu$. Thus, a number different from all numbers in the sequence containing all rational numbers is created. Therefore, it must be an irrational number. This will be the first irrational number generated and will be incorporated into the sequence $S_{r}$ as the first element, which shifts all elements in the original sequence from the $n$-th to the $n+1$-th place. This recursive construction will be repeated; it will be applied again on the newly generated set $S_{r}$, generating in the same way the second irrational number and placing it in the sequence $S_{r}$ in the first position, shifting all elements in $S_{r}$ again, then the third irrational number, and so on. All numbers generated in this way are well defined, all digits $d_{\nu\mu}$ of such generated irrational numbers $D_{1},D_{2},\ldots,D_{\nu},\ldots$ are defined and known. This is because all digits in the initial sequence $S_{r}$ that contains all rational reals are defined. In creating irrational numbers by Cantor’s diagonal construction, there is no freedom because each diagonal element 0 must be switched to 1, or 1 to 0; one must apply $d_{\nu\nu}\neq a_{\nu\nu}$ for each diagonal element. Since in the initial sequence $S_{r}$ of rational numbers all $a_{\nu\mu}$ are known and fixed, all numbers $D_{1}$, $D_{2}$, …that are produced with the diagonal construction are defined.
@@
To generate the first layer in our model, the construction will be modified so that in the first step, not only one but infinitely many irrational numbers are generated. This is possible because all decimal places $a_{11},a_{12},\ldots,a_{1\mu},\ldots$ can be used simultaneously and an infinite number of diagonal constructions can be applied simultaneously in the first step. There will be a diagonal construction that starts with the first decimal place $a_{11}$, one that starts with the second decimal place $a_{12}$, one with the third $a_{13}$, and so on. New irrational numbers will be created using an infinite number of diagonal constructions, each starting from a different decimal place. After step one, the first layer, denoted as $\Sigma_{1}$, will consist of an infinite sequence of $\aleph_{0}$ rational numbers from the 00-th layer and an infinite sequence of $\aleph_{0}$ irrational numbers produced by an infinite number of diagonal constructions applied to the set of rational reals of the 00-th layer. These two sets, the 00-th layer and the irrational numbers generated by diagonal constructions, will be combined into one sequence by alternating the first number from the 00-th layer with the first number from the list of created irrational numbers, and so on, to create the first layer, the $\Sigma_{1}$ set. The construction from step one will be repeated, applying an infinite number of diagonal constructions to the sequence that represents the just created set $\Sigma_{1}$, creating an additional $\aleph_{0}$ irrational numbers that will then be combined with set $\Sigma_{1}$ to create set $\Sigma_{2}$. This recursive construction will continue indefinitely, creating an additional $\aleph_{0}$ of newly created irrational numbers each time.
@@
However, this construction will conclude after $\omega$ steps due to the atomistic structure of real numbers, which prevents endless division of an interval. The process concludes exactly after $\omega$ divisions because an infinite division of intervals results in singletons that cannot be divided further. At each step of creating layers, infinite amounts of irrational numbers are inserted between the algebraic numbers of the 00-th layer, dividing each segment defined by a pair of rational numbers further until it becomes a singleton, as follows from the BW theorem or the NIP.
Proof.
This Theorem is equivalent to Theorem 1. In both cases, irrational numbers are generated and inserted between rational numbers, subdividing the intervals into smaller and smaller segments. Decimal places $a_{\nu\mu}$ that represent numbers in (15) are equivalent to nested intervals in $S_{Q}$, and making chnages in decimal places in (15) is equivalent to marking numbers in intervals $S_{Q}$. In both cases generated is a set of everywhere dense irrational numbers. While in Theorem 1 completnes of set $S_{m}$ arrises from its density and the fact that any number must be marked, and that there is no exclusion of numbers, here completeness follows from the fact that any missing number generated by diagonal argument is actullay part of the set of irrational numbers generated by this construction. There is also a proof that any irrational number can be generated by a diagonal sequence $D_{1},D_{2},\ldots,D_{\nu},\ldots$ in the set of rational numbers, by Gray, theorem 3 in [22]. He proved that by making permutations in the positions of the elements of the initial rational sequence it is possible to obtain the condition that will generate the specific targeted irrational number. By our construction, we are not creating direct permutations of the positions of the elements in the rational sequence, but instead doing that indirectly by making in each step an infinite number of changes of the elements in the initial sequence, by adding to the initial sequence in each step an infinite number of new irrational numbers that have decimal places different from the initial sequence. Doing that using an infinite number of diagonals, changing all digits on each of the diagonals, we are actually in each step creating an infinite number of permuted initial sets considered by Gray. The permutations we are creating are predetermined by the initial set, but the fact that generated irrational numbers have a random distribution of the digits, ensures that conditions for generating any number can be achieved, i.e., that method does not exclude any number that one can specify. Any number that one can specify can be in principle generated, because the construction is going indefinitely, it is only a question in which step it will be produced. Actually, if one specifies all decimals of a number, it can be calculated at which sep that number will be generated.
@@
The diagonal argument cannot be used to create missing numbers not generated by this model because it is utilized to generate new numbers. Any ‘missing’number produced by diagonal construction is, in fact, just a number that belongs to this model. ∎
@@
5 Remarks to the non-denumerability proofs
There are many variations of proofs of the uncountability of real numbers. However, fundamentally, all of them share a common basis: they revolve around infinite nested sequences of intervals. We will discuss the critical point that appears in all these proofs, the interpretation of the intersection of these intervals and the nature of the number that resides within the intersection. According to the NIP, the intersection of an infinite nested sequence of closed nonempty intervals must be a zero-size interval, a singleton, or a single number. This number must be one of the elements from the intervals of the sequence, which, in all uncountability proofs, are the numbers of the supposed denumerable sequence assumed to include all numbers. Contrarily, in all nondenumerability proofs, that number is declared as the number that is outside the denumerable sequence, and its existence is used as the argument that the sequence does not contain all numbers. This discrepancy between the NIP, which posits that the number is a member of the sequence, and the uncountability proofs, which assert the opposite, will be elaborated upon, considering the proofs from 1873 and 1879, and the diagonal argument.
@@ 5 Remarks to the non-denumerability proofs
In the concluding arguments of both proofs, two fundamental issues arise: the technical problem that the proposed construction of a number different from all listed numbers cannot be completed, and the logical problem that assumes if a number does not have an assigned specific $n$-th place in the sequence, it is not in the sequence. It will be demonstrated that the number proposed to be different from all others cannot be constructed as intended. What is actually constructed is the number $\eta$, which is not distinct from other numbers. The construction requires an additional step to differentiate it from the listed numbers, but as it will be shown this final step cannot be completed as proposed. Also, the constructed number $\eta$ does not have a specific $n$-th place in the sequence due to the method of its construction; however, this does not imply that it is not part of the sequence. Let us elaborate on both of these points.
@@
The claim is that in the limit as $\nu\to\infty$, where $\omega_{a}^{\nu}=A$ and $\omega_{b}^{\nu}=B$, and when $A=B$, there exists an open interval $(A,B)$ that eliminates all endpoints $\omega_{a}^{\nu}$, $\omega_{b}^{\nu}$, and is not empty, containing the number $\eta=A=B$. To quote: “because $\eta$ is in the interior of the interval $[\omega_{a}^{\nu},\omega_{b}^{\nu}]$, but $\omega_{\nu}$ lies outside of it. For all $\nu,\eta\in(\omega_{a}^{\nu},\omega_{b}^{\nu})$ but $\omega_{\nu}\notin(\omega_{a}^{\nu},\omega_{b}^{\nu})$.”
@@
The numbers $\eta$ and $\omega_{n}$ are then defined through an infinite sequence of nested intervals
@@
In the nested interval representation, the equivalent expression stipulates that at each finite step $n$, the process involves considering the intersection (20) of the first $n$ intervals $I_{n}$ representing the first $n$ decimals of the numbers $\eta$, and $\omega_{n}$. Within these intersections, distinct subintervals $H_{n}$ for $\eta$ and $\Omega_{n}$ for $\omega_{n}$ are selected to ensure divergence. Thus the divergence is created by choosing disjunctive subintervals to guarantee that $\eta$ and $\omega_{n}$ differ at the $n$-th stage, effectively making the numbers different at each $n$-th decimal place or $n$-th interval, because they will have different intersections (20):
@@
The diagonal argument assumes that the process of creating disjunctive subintervals $H_{i}$ and $\Omega_{i}$ can continue indefinitely, theoretically allowing for the construction of the number $\eta$ different from all listed numbers $\omega_{n}$. However, this seemingly straightforward assumption contradicts the NIP. At each finite stage in sequence (19), every interval $I_{n}=[\alpha_{n},\beta_{n}]$ has a finite size, with $\alpha_{n}<\beta_{n}$ consistently satisfied. This ensures that each interval $I_{n}$ contains at least two disjunctive subintervals $H_{n}$ and $\Omega_{n}$, necessary to construct number $\eta$ such that it will be different from listed numbers $\omega_{n}$. However, this intuitively natural requirement cannot be met when applied to the entire sequence (19) as its intersection interval (20), in the limit, has zero size.
In the limit, as $n\rightarrow\infty$, $\alpha_{n}\rightarrow A$ and $\beta_{n}\rightarrow B$, and $A=B$, the intersection (20) of all intervals $I_{n}=[\alpha_{n},\beta_{n}]$, i.e., the intersection interval $[A,B]$, has zero size and contains only a single number. So, there is no possibility that in the intersection interval will exist two disjunctive subintervals $H_{n}$ and $\Omega_{n}$, which is required by diagonal argument to make the intersection (21) representing number $\eta$ different from the intersections (22) representing a listed number $\omega_{n}$.
The intersection theorem dictates that sequence (19) has a zero-size intersection with only one number in it. Intervals $H_{n}$ representing decimals $a_{n}$ in number $\eta$ and intervals $\Omega_{n}$ representing $\omega_{nn}$ in $\omega_{n}$ for every $n$ belong to the same interval $I_{n}=[\alpha_{n},\beta_{n}]$ because they represent the same $n$-th decimal position. Therefore, they must lie within the same intersection (20) of intervals $I_{n}$ at each step for any $n$. For every $n$ in the construction of numbers $\eta$ and $\omega_{n}$, the intersection of the first $n$ intervals $[\alpha_{n},\beta_{n}]$ has non-zero size and contains disjunctive subintervals $H_{n}$ and $\Omega_{n}$, allowing $\eta$ to differ from the first $n$ listed numbers $\omega_{n}$ in their $n$-th decimal places. However, at these finite stages, number $\eta$ is not yet defined or constructed, regardless of how large $n$ is. Number $\eta$ is only defined or constructed in the limit as $n\rightarrow\infty$, when $\alpha_{n}\rightarrow A$ and $\beta_{n}\rightarrow B$. But in this limit, the intersection $[A,B]$ is a singleton containing only one number. The requirement at each $n$-th step to select two disjunctive subintervals $H_{n}\neq\Omega_{n}$ that make sequence intersections (21) and (22) different, cannot be fulfilled in the limit because the intersection (20) becomes a singleton, a single number. So, the number $\eta$, that is different from all listed numbers cannot be constructed since in the limit do not exist two disjunctive subintervals, required to make the number $\eta$ different from all listed $\omega_{n}$ numbers.